OpenStudy (anonymous):

Rachel moves a magnet toward a metal paperclip. What will most likely happen if she continues moving it toward the paperclip?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@EclipsedStar

OpenStudy (anonymous):

@sleepyjess

OpenStudy (ybarrap):

For a bar magnet, the magnetic field (far away) is approximately $$ \bf{B}={\mu_om\over 2\pi z^3},\text{ where } m=M_0\pi a^2 l $$ \(z \) is the distance from the magnet, \(a \) is the radius, \(l \) is the length of the magnet and \(M_0\) is the magnetic dipole moment density. The important point is that the magnetic force falls off as \(\large {1\over z^3}\). So the closer she gets to the paper clip, the stronger the magnetic force: